OpenStudy (anonymous):

Rewrite with a rational exponent 4 square root 21 looks like a square root division sign the 4 is kind of powered and 21 is under the medium long part

OpenStudy (anonymous):

\[\sqrt[4]{21}\]

OpenStudy (campbell_st):

\[\sqrt[4]{21} = 21^{\frac{1}{4}}\]

OpenStudy (anonymous):

21^1/4

OpenStudy (anonymous):

ops the same ur ansewer campbell

OpenStudy (anonymous):

thanks cause I got 21^1/4 but wanted to know i was on the right path

OpenStudy (anonymous):

good luck lylas